{3x - y=2
{-2 + 5y=6

Answer by RAY100(1637)

(1) 3x -y =2,,,,( 3x -2) =y
.
(2) -2x +5y =6,,,,,,,,,,,,,,added x to complete the problem
.
subst (1) into (2)
.
-2x + 5 (3x -2) = 6
.
-2x +15x -10 =6
.
13x = 16
.
x=16/13
.
subst in (1),,,,y = 3 (16/13) -2 = (48 - 26) / 13 = 22/13
.
checking
.
3(16/13) -(22/13) = (48 - 22) /13 = 26 / 13 = 2,,,,ok
.
-2(16/13) + 5 ( 22/13) = ( -32 + 110) / 13 = 78 /13 =6,,,,,ok